OpenStudy (anonymous):

find the equation of a circle whose endpoints of a diameter are P(-1,1) and Q(5,9)

OpenStudy (jchick):

Standard form of equation for a circle:(x-h)^2 + (y - k)^2 = r^2

OpenStudy (jchick):

(h,k)=(x,y) coordinates of center, r=radius

OpenStudy (jchick):

Use distance formula to find radius: diameter=√(x1-x2)^2+(y1-y2)^2)=√((5+1)^2+(9-1)^2)=√36+64

OpenStudy (jchick):

=√100=10 radius=10/2=5

OpenStudy (jchick):

coordinates of center=midpoint of given points: (-1,1) and (5,9) midpoint=(x1+x2)/2,(y1+y2)/2=(-1+5)/2),(9+1)/2)=(2,5) center: (2,5)

OpenStudy (jchick):

Equation of given circle:

OpenStudy (jchick):

(x - 2)^2 + (y - 5)^2 = ??

OpenStudy (jchick):

25